OpenStudy (anonymous):

How do you simplify (x-8)^2

OpenStudy (anonymous):

(a - b)^2 = a^2 - 2ab + b^2. given that a = x and b = 8, can you solve it from there?

OpenStudy (anonymous):

you don't you can however "multiply" start with \[(x-8)(x-8)\] then multiply it out

OpenStudy (anonymous):

We meet again, satellite! ☺

OpenStudy (anonymous):

\[x ^{2} - 2*x*8 + 8^{2}\]

OpenStudy (anonymous):

That simplifies to \[x^2 -16x + 64\]. and there you go.